OpenStudy (anonymous):

2 and 5/6 * 8 and 4/7= ?? Answer needs to be a mixed number in simplest terms

OpenStudy (anonymous):

are the "and"s multiplication?

OpenStudy (anonymous):

no they are mixed numbers

OpenStudy (anonymous):

so: \[\large 2\frac{ 5 }{ 6 } \times 8\frac{ 4 }{ 7 }\]

OpenStudy (anonymous):